# An introduction to reinforcement learning for neuroscience

### v2

@@ 8 Distributional reinforcement learning
  • When independently fitting a slope to the data above ($\alpha_{i}^{+}$) and below ($\alpha_{i}^{-}$) the reversal point of neuron $i$, the reversal point of the neuron was positively correlated with $\tau_{i}=\frac{\alpha_{i}^{+}}{\alpha_{i}^{-}+\alpha_{i}^{+}}$. This is consistent with the expectile distributional RL setting, where the reversal point is $V_{\tau_{i}}\approx\epsilon_{\tau_{i}}$ (Figure 6C).
@@
These findings generalize the canonical RPE view of Schultz et al., (1997), which can be seen as the averaged version of the theory by Dabney et al., (2020). The expectile regression algorithm investigated by Dabney et al., (2020) relies on non-local TD updates and non-linear ‘imputation’ of the return distribution $p_{\{V_{\tau_{i}}\}}(\tilde{Z}^{\pi}(s))$ induced by the learned expectiles $\{V_{\tau_{i}}(s)\}$. However, recent work has suggested more biologically plausible distributional RL updates to address these challenges (Tano et al., , 2020; Wenliang et al., , 2023).
@@ 9 Learning from scalar rewards?
The methods discussed so far have all relied on some scalar ‘reward’ available from the environment. Distributional RL predicted the full distribution of possible rewards but still assumed that an external reward was eventually observed. This is perhaps reasonable in many experimentally controlled settings, where we deliver juice or water to animals as a function of their actions. It is also appropriate in machine learning, where we want to optimize some externally imposed objective that can be quantified. However, in natural environments, there is no such natural definition of ‘reward’ – instead, different actions can be rewarded in different ways that must be converted to an internal learning signal (Juechems and Summerfield, , 2019). For example, foraging for food might reduce hunger, while collecting water might reduce thirst. Starting a family might have an initial energy cost, but it could improve survival later in life and propagate the gene pool. An RL purist would convert all of these different gains and losses into a single common unit and balance them appropriately to decide what actions to take (Silver et al., , 2021). However, it has also been suggested that biological agents adaptively change their instantaneous objective via a higher-order controller in e.g. prefrontal cortex, which determines what the current lower-level objective is (Juechems and Summerfield, , 2019; Miller and Cohen, , 2001; Botvinick, , 2008). This resembles ideas in hierarchical reinforcement learning (Pateria et al., , 2021; Botvinick, , 2008), but here the higher-order policies are not necessarily learned through reinforcement learning within a lifetime, instead emerging during evolution where survival depends on the discovery of ‘useful’ objectives.
@@
A related challenge is the technical difficulty of learning from scalar rewards, which are generally sparse. While the methods in Section 7 can train a neural network to maximize reward in theory, in practice they are often noisy, unstable, or find local minima. As a topical example, consider the challenge of training large language models to interact with human users. Training such a model from scratch using reinforcement learning would be near-impossible, but ‘pretraining’ the model on a large unsupervised dataset followed by ‘reinforcement learning from human feedback’ has proven hugely successful and revolutionized language models for human interactions (Team et al., , 2023). This works because the large unlabelled dataset provides a way to learn good representations that distinguish different concepts, since this is necessary to solve the base task of predicting the next word. Once such representations are learned, the subsequent finetuning for human interaction is an easier problem that can be solved with reinforcement learning.
@@
Of more relevance to neuroscience, it is also common in deep RL to introduce additional ‘auxiliary costs’ to the utility function that are jointly optimized together with reward maximization, and which can use a richer data source to drive representation learning in the network (Jaderberg et al., , 2016). A popular approach is to include losses that require the agent to predict the next observation from the current state and action (Jaderberg et al., , 2016; Zintgraf et al., , 2019). Such predictive losses have close parallels in neuroscience, where it has been suggested that predictive learning could drive many of the representations observed in biological circuits (Rao and Ballard, , 1999; Stachenfeld et al., , 2017; Whittington et al., , 2020; Blanco-Pozo et al., , 2021) and serve as a foundation for model-based planning (Jensen et al., , 2023). Fang and Stachenfeld, (2023) also recently showed that augmenting an RL agent with an auxiliary predictive objective leads to neural representations that resemble biological circuits more closely. This suggests a potentially important interaction between self-supervised representation learning and reward-driven reinforcement learning in biological circuits.
@@
If a good predictive model of the world is learned, this model can also be used together with a search algorithm to turn the reinforcement learning problem into a supervised learning problem. As an example, the MuZero model developed by Schrittwieser et al., (2020) to learn Atari, chess, shogi, and Go, was trained to predict future values and actions by unrolling the environment using a learned latent representation. It then used Monte Carlo tree search (MCTS; a model based-search algorithm c.f. Section 5) and the predicted value function to improve the provisional actions predicted by the base policy network. These final actions, which had been optimized using MCTS, were treated as a supervised learning target for the base policy network. Through many iterations of such predictive learning, MCTS-based policy improvement, and training of the base policy, the network learned both the transition structure of the task and to select good actions. Such semi-supervised approaches to reinforcement learning are useful in the common setting where learning the transition structure of the world is easier than learning a policy (Jensen et al., , 2023). These methods also have interesting parallels to learning in biological networks, where interactions between model-based and model-free systems are similarly thought to drive action selection based on learned latent representations (Botvinick et al., , 2020).
@@
Finally, algorithms that seek to maximize scalar rewards often run into challenges related to exploration. In particular, once an above-average policy has been identified, exploration is disfavored because it has lower expected reward, even if there is a finite probability of discovering new and better policies. This is why imposing stochasticity in the form of e.g. $\epsilon$-greedy policies is common in the tabular RL literature. In deep RL, it is instead popular to include various forms of ‘exploration bonuses’ in the objective function. Policy gradient algorithms for example often add an auxiliary entropy loss of the form $\mathcal{L}_{E}=\sum_{\tau\sim p_{\pi_{\theta}}(\tau)}\sum_{t=0}^{T}\sum_{a}\pi_{\theta}(a_{t}|s_{t})\log\pi_{\theta}(a_{t}|s_{t})$, and indeed the simple example in Figure 5 does not work without such an entropy loss. Other approaches to improve exploration include hierarchical reinforcement learning, which introduces autocorrelations in the policy to improve coverage of the state space. This is reminiscent of the biological ‘Lévy flight’ hypothesis, which suggests that animals explore an environment using a heavy-tailed distribution of ‘step sizes’ to maximize the probability of finding sparse rewards (Viswanathan et al., , 1999). Biological agents are also thought to engage in periods of ‘random exploration’ that can be interpreted as a biological parallel to $\epsilon$-greedy-like algorithms. Random exploration appears to be under noradrenergic control (Tervo et al., , 2014; Dubois et al., , 2021), while the zona incerta can drive more directed exploration (Ahmadlou et al., , 2021). Such neural control of exploration is particularly well-characterized in the zebra finch song circuit (Ölveczky et al., , 2005), and it is consistent with a view of top-down imposition of flexible reward functions – where one possible objective could be to increase variability or reduce uncertainty – rather than optimization of a global scalar reward.
@@ 10 Discussion
In this review, we have provided a mathematical overview of some of the many reinforcement learning methods that are commonly used in systems and computational neuroscience. We have also highlighted a range of explicit parallels between these methods and experimental results in neuroscience and cognitive science to illustrate the utility of reinforcement learning as a framework for understanding biological learning and decision making. This has ranged from classical work on reward prediction errors (Schultz et al., , 1997) to recent findings of distributional reinforcement learning in biological circuits (Dabney et al., , 2020).
@@
While RL has already had a profound influence on systems neuroscience, several open questions remain. In particular, much work has focused on simple stimulus-response or binary decision making tasks. This is a far reach from ethologically relevant problems that involve processing multimodal stimuli, decision making with long-lasting consequences, and high-dimensional motor control. Some recent work building on deep RL has started to bridge this gap. For example, Banino et al., (2018) demonstrated the emergence of grid cells in agents navigating complex environments, Aldarondo et al., (2024) showed similarities between an RL agent trained to control a virtual rodent and corresponding biological motor representations, and Jensen et al., (2023) showed parallels between a recurrent meta-RL agent and human behaviour in a navigation task requiring temporally extended thinking. However, much work remains to extend our neuroscientific understanding to ethologically relevant settings, both experimental and computational.
@@
A related challenge will be to combine different components of existing models to capture the generalist nature of biological circuits. This is in contrast to past work, which has often focused on a single neural circuit or function, such as motor control or navigation. Such a generalist approach will involve explicit modeling of the roles of different brain regions, and more importantly it will require us to capture how they interact with one another during learning and decision making. Clearly, such models will need to be constrained by experimental data, both at the level of behaviour and at the level of neural activity. This is becoming increasingly feasible with recent advances in recording technologies, both for high-resolution behavioural tracking (Mathis et al., , 2018; Dunn et al., , 2021) and for simultaneous and long-term recording of neural activity (Steinmetz et al., , 2021; Pachitariu et al., , 2016; Dhawale et al., , 2017).
@@
Finally, most work on reinforcement learning in neuroscience has considered short-term decision making tasks, where planning and decision making in primitive state and action spaces are feasible. This is in stark contrast to most human decision making, which occurs over extended timescales and often involves hierarchies of decisions or priorities that change over time. For example, we may decide to pursue an undergraduate degree at Cambridge University, which then requires us to (i) write an application, (ii) prepare for an interview, and (iii) arrange our travel. Each of these processes in turn require us to plan increasingly low-level decisions, such as booking a flight or deciding which bus to take to the airport. This is the topic of hierarchical reinforcement learning, which has already been highlighted as a potentially useful model of human behaviour (Eckstein and Collins, , 2020; Botvinick, , 2008; Botvinick et al., , 2009) and is becoming an increasingly important area of research in machine learning (Pateria et al., , 2021). While reinforcement learning will undoubtedly remain important for understanding such flexible animal behaviour, we should also keep in mind the shortcomings and challenges of learning from scalar rewards (Section 9). It will be important to investigate the regimes in which reinforcement learning provides a good model of behaviour and neural dynamics, while also exploring other frameworks with potentially richer learning substrates, and studying the interactions between such different learning algorithms.

### v3

@@ 8 Distributional reinforcement learning
  • When independently fitting a slope to the data above ($\alpha_{i}^{+}$) and below ($\alpha_{i}^{-}$) the reversal point of neuron $i$, the reversal point of the neuron was positively correlated with $\tau_{i}=\frac{\alpha_{i}^{+}}{\alpha_{i}^{-}+\alpha_{i}^{+}}$. This is consistent with the expectile distributional RL setting, where the reversal point is $V_{\tau_{i}}\approx\epsilon_{\tau_{i}}$, which is monotonic in $\tau_{i}$ (Figure 6C).
@@
These findings generalize the canonical RPE view of Schultz et al. [97], which can be seen as the averaged version of the theory by Dabney et al. [20]. The expectile regression algorithm investigated by Dabney et al. [20] relies on non-local TD updates and non-linear ‘imputation’ of the return distribution $p_{\{V_{\tau_{i}}\}}(\tilde{Z}^{\pi}(s))$ induced by the learned expectiles $\{V_{\tau_{i}}(s)\}$. However, recent work has suggested more biologically plausible distributional RL updates to address these challenges [111, 128].
@@ 9 Learning from scalar rewards?
The methods discussed so far have all relied on some scalar ‘reward’ available from the environment. Distributional RL predicted the full distribution of possible rewards but still assumed that an external reward was eventually observed. This is perhaps reasonable in many experimentally controlled settings, where we deliver juice or water to animals as a function of their actions. It is also often appropriate in machine learning, where we want to optimize some externally imposed objective that can be quantified. However, in natural environments, there is no such ‘global’ reward – instead, different actions can be rewarded in different ways that must be converted to an internal learning signal [49]. For example, foraging for food might reduce hunger, while collecting water might reduce thirst. Starting a family might have an initial energy cost, but it could improve survival later in life and propagate the gene pool. An RL purist would convert all of these different gains and losses into a single common unit and balance them appropriately to decide what actions to take [99]. However, it has also been suggested that biological agents adaptively change their instantaneous objective via a higher-order controller in e.g. prefrontal cortex, which determines what the current lower-level objective is [49, 67, 15]. This resembles ideas in hierarchical reinforcement learning [85, 15], but here the higher-order policies are not necessarily learned through reinforcement learning within a lifetime, instead emerging during evolution where survival depends on the discovery of ‘useful’ objectives.
@@
A related challenge is the technical difficulty of learning from scalar rewards, which are generally sparse. While the methods in Section 7 can train a neural network to maximize reward in theory, in practice they are often noisy, unstable, or find local minima. As a topical example, consider the challenge of training large language models to interact with human users. Training such a model from scratch using reinforcement learning would be near-impossible, but ‘pretraining’ the model on a large unsupervised dataset followed by ‘reinforcement learning from human feedback’ has proven hugely successful and revolutionized language models for human interactions [112]. This works because the large unlabelled dataset provides a way to learn good representations that distinguish different concepts, since this is necessary to solve the base task of predicting the next word. Once such representations are learned, the subsequent finetuning for human interaction is an easier problem that can be solved with reinforcement learning.
@@
Of more relevance to neuroscience, it is also common in deep RL to introduce additional ‘auxiliary costs’ to the utility function that are jointly optimized together with reward maximization, and which can use a richer data source to drive representation learning in the network [46]. A popular approach is to include losses that require the agent to predict the next observation from the current state and action [46, 134]. Such predictive losses have close parallels in neuroscience, where it has been suggested that predictive learning could drive many of the representations observed in biological circuits [89, 103, 129, 11] and serve as a foundation for model-based planning [47]. Fang and Stachenfeld [33] also recently showed that augmenting an RL agent with an auxiliary predictive objective leads to neural representations that resemble biological circuits more closely. This suggests a potentially important interaction between self-supervised representation learning and reward-driven reinforcement learning in biological circuits.
@@
If a good predictive model of the world is learned, this model can also be used together with a search algorithm to turn the reinforcement learning problem into a supervised learning problem. As an example, the MuZero model developed by Schrittwieser et al. [95] to learn Atari, chess, shogi, and Go, was trained to predict future values and actions by unrolling the environment using a learned latent representation. It then used Monte Carlo tree search (MCTS; a model based-search algorithm c.f. Section 5) and the predicted value function to improve the provisional actions predicted by the base policy network. These final actions, which had been optimized using MCTS, were treated as a supervised learning target for the base policy network. Through many iterations of such predictive learning, MCTS-based policy improvement, and training of the base policy, the network learned both the transition structure of the task and to select good actions. Such semi-supervised approaches to reinforcement learning are useful in the common setting where learning the transition structure of the world is easier than learning a policy [47]. These methods also have interesting parallels to learning in biological networks, where interactions between model-based and model-free systems are similarly thought to drive action selection based on learned latent representations [14].
@@
Finally, algorithms that seek to maximize scalar rewards often run into challenges related to exploration. In particular, once an above-average policy has been identified, exploration is disfavored because it has lower expected reward, even if there is a finite probability of discovering new and better policies. This is why imposing stochasticity in the form of e.g. $\epsilon$-greedy policies is common in the tabular RL literature. In deep RL, it is instead popular to include various forms of ‘exploration bonuses’ in the objective function. Policy gradient algorithms for example often add an auxiliary entropy loss of the form $\mathcal{L}_{E}=\sum_{\tau\sim p_{\pi_{\theta}}(\tau)}\sum_{t=0}^{T}\sum_{a}\pi_{\theta}(a_{t}|s_{t})\log\pi_{\theta}(a_{t}|s_{t})$, and indeed the bandit example in Figure 5 does not work without such an entropy loss. Other approaches to improve exploration include hierarchical reinforcement learning, which introduces autocorrelations in the policy to improve coverage of the state space. This is reminiscent of the biological ‘Lévy flight’ hypothesis, which suggests that animals explore an environment using a heavy-tailed distribution of ‘step sizes’ to maximize the probability of finding sparse rewards [122]. Biological agents are also thought to engage in periods of ‘random exploration’ that can be interpreted as a biological parallel to $\epsilon$-greedy-like algorithms. Random exploration appears to be under noradrenergic control [113, 29], while the zona incerta can drive more directed exploration [2]. Such neural control of exploration is particularly well-characterized in the zebra finch song circuit [81], and it is consistent with a view of top-down imposition of flexible reward functions – where one possible objective could be to increase variability or reduce uncertainty – rather than optimization of a global scalar reward.
@@ 10 Discussion
In this review, we have provided a mathematical overview of some of the many reinforcement learning methods that are commonly used in systems and computational neuroscience. We have also highlighted a range of explicit parallels between these methods and experimental results in neuroscience and cognitive science to illustrate the utility of reinforcement learning as a framework for understanding biological learning and decision making. This has ranged from classical work on reward prediction errors [97] to recent findings of distributional reinforcement learning in biological circuits [20].
@@
While RL has already had a profound influence on systems neuroscience, several open questions remain. In particular, much work has focused on simple stimulus-response or binary decision making tasks. This is a far reach from ethologically relevant problems that involve processing multimodal stimuli, decision making with long-lasting consequences, and high-dimensional motor control. Some recent work building on deep RL has started to bridge this gap. For example, Banino et al. [6] demonstrated the emergence of grid cells in agents navigating complex environments, Aldarondo et al. [4] showed similarities between an RL agent trained to control a virtual rodent and corresponding biological motor representations, and Jensen et al. [47] showed parallels between a recurrent meta-RL agent and human behaviour in a navigation task requiring temporally extended thinking. However, much work remains to extend our neuroscientific understanding to ethologically relevant settings, both experimental and computational.
@@
A related challenge will be to combine different components of existing models to capture the generalist nature of biological circuits. This is in contrast to past work, which has often focused on a single neural circuit or function, such as motor control or navigation. Such a generalist approach will involve explicit modeling of the roles of different brain regions, and more importantly it will require us to capture how they interact with one another during learning and decision making. Clearly, such models will need to be constrained by experimental data, both at the level of behaviour and at the level of neural activity. This is becoming increasingly feasible with recent advances in recording technologies, both for high-resolution behavioural tracking [63, 30] and for simultaneous and long-term recording of neural activity [105, 82, 27].
@@
Finally, most work on reinforcement learning in neuroscience has considered short-term decision making tasks, where planning and decision making in primitive state and action spaces are feasible. This is in stark contrast to most human decision making, which occurs over extended timescales and often involves hierarchies of decisions or priorities that change over time. For example, we may decide to pursue an undergraduate degree at Cambridge University, which then requires us to (i) write an application, (ii) prepare for an interview, and (iii) arrange our travel. Each of these processes in turn require us to plan increasingly low-level decisions, such as booking a flight or deciding which bus to take to the airport. This is the topic of hierarchical reinforcement learning, which has already been highlighted as a potentially useful model of human behaviour [31, 15, 17] and is becoming an increasingly important area of research in machine learning [85]. While reinforcement learning will undoubtedly remain important for understanding such flexible animal behaviour, we should also keep in mind the shortcomings and challenges of learning from scalar rewards (Section 9). It will be important to investigate the regimes in which reinforcement learning provides a good model of behaviour and neural dynamics, while also exploring other frameworks with potentially richer learning substrates. This will also enable the study of interactions between such different learning algorithms, which will likely be necessary to understand biological learning in rich environments.
Acknowledgements
Guillaume Hennequin, Will Dorrell, and the NBDT reviewers and editor provided valuable feedback that improved and clarified the manuscript. Most of the content in this review was originally prepared for the 2023 Janelia Theoretical Neuroscience Workshop.